Soybean meal is 16% protein; cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320-lb. mixture with 15% protein?
How many pounds of the cornmeal should be in the cornmeal mixture?
How many pounds of the soybean meal should be in the mixture?
HELP I hate word problems!!!! PLEASE!

Answer by scott8148(6628)

let x=soybean, so 320-x=corn

x(16%)+(320-x)(8%)=320(15%) __ multiplying by 100 to clear percentages __ 16x+2560-8x=4800

subtracting 2560 __ 8x=2240 __ dividing by 8 __ x=280

substituting __ 320-(280)=40